A standard teaspoon contains about 5 ml of liquid. If a single administration of castor oil consists of 15 ml, with a maximum number of four administrations per day, how many teaspoons of castor oil are there in a day’s dose?

A dose of methylnaltrexone bromide is administered to patients weighing over 251 lbs at a level of 0.15 mg per kilogram of body mass. Each dose is administered through a subcutaneous injection every other day. If a patient weighs 264 lbs and the methylnaltrexone bromide solution has a concentration of 20 mg/ml, how many milliliters will be administered in a period of two weeks?
Answer #1

ANSWER.

* Calculation of amount of teaspoons of castor oil required in a day.

          Amount of castor oil present in a single administration = 15 ml.

           Number of administrations per day = 4

         Amount of castor oil required per day = 15 * 4

                                                                             =60 ml.

             1 teaspoon = 5 ml.

      So

        Number of teaspoons of castor oil needed per day = 60 / 5

                                                                                          = 12 teaspoons.

* Calculation of amount(ml) of drug needed.

         Formula = Desired dose/ Dose in hand * Quantity.

               Desired dose(D) = 0.15 mg/kg.

               Weight of the patient = 264 lbs.

            Convert lbs value to kg for calculation

             So

                      1 kg = 2.205 lbs.

               Weight of the patient = 264 /2.205

                                                 = 119.7 kg.

                                                 = 120 kg.

                Desired dose for a patient with 120 kg = 0.15 * 120

                                                            = 18 mg.

                    Dose in hand (H) = 20 mg.

                           Quantity (Q) = 1 ml.

         As per formula

                 Amount (ml) of drug required = D/H *Q

                                                                = 18 / 20 * 1

                                                                = 0.9 * 1

                                                               = 0.9 ml.